Americium-241 is used in some smoke detectors. It is an alpha emitter with a half-life of 432 years. How long will it take in years for 44.0 % of an Am-241 sample to decay?
Answer #1

t1/2 = 0.693/K

K = 0.693 / 432

K = 1.6042 * 10^-3

K = 2.303/t log(a / (a-x))

1.6042 * 10^-3 = 2.303 / t log (100/(100-44))

t = 361.5 years.